# LAMBench: A Benchmark for Large Atomic Models

### v1

@@ II.2 Benchmark results
Table 2 presents the LAMBench Leaderboard, showcasing the performance of LAMs in terms of their generalizability and applicability. Currently, adaptability is excluded from the comparison due to the absence of this test for most of the LAMs under evaluation. This leaderboard aims to strike a balanced comparison between key performance aspects and is intended to offer insights into model suitability for applications relevant to real-world scientific discovery.
@@
DPA-2.4-7M 0.265 0.208 0.614 0.039
GRACE-2L-OAM 0.340 0.262 0.678 0.309
SevenNet-l3i5 0.355 0.240 0.279 0.036
MACE-MPA-0 0.356 0.291 0.291 0.000
Orb-v2 0.356 0.560 1.343 2.649
SevenNet-MF-ompa 0.358 0.300 0.088 0.000
MatterSim-v1-5M 0.389 0.280 0.388 0.000
MACE-MP-0 0.405 0.341 0.291 0.089
@@
Figure 2: Dimensionless error metrics for generalizability tasks across different domains. (a) Dimensionless error metrics for force field tasks across five distinct domains. (b) Property-calculation tasks: the Small Molecules domain is assessed using the TorsionNet500 benchmark, while the Inorganic Materials domain is evaluated using the phonon MDR benchmark. To reassess DPA-2.4-7M with enhanced XC functional alignment, referred to as DPA-2.4-7M-bestXC, we selected task heads trained on Yang2023ab for the ANI-1x and AIMD-Chig datasets. For the TorsionNet500 benchmark, we adopted the task head trained on the SPICE2 dataset. For the Sours2023Applications dataset, we applied a D3(BJ) dispersion correction.
@@
In the force field generalizability task, as illustrated in Table 2, performance is evaluated using the dimensionless error $\bar{M}_{\mathrm{FF}}^{m}$. Notably, DPA-2.4-7M demonstrates substantially greater generalizability compared to other LAMs. Interestingly, aside from MatterSim-v1-5M and MACE-MP-0, most models exhibit only minor differences in generalizability. To further investigate the generalizability of these LAMs across specific domains, we present the domain-wise dimensionless error metric in Figure 2(a). In general, most LAMs exhibit relatively low inference errors in the domains of Inorganic Materials, Reactions, and Small Molecules, while showing inferior performance in the domains of Catalysis and Biomolecules and Supramolecules. In the Reactions domain, DPA-2.4-7M, Orb-v2, and SevenNet-MF-ompa achieve the lowest errors, with MatterSim-v1-5M following closely behind. For Small Molecules, DPA-2.4-7M and Orb-v2 rank first and second, respectively, outperforming all other models in this domain. In the Inorganic Materials domain, the comparative performance of the models shows significant alignment with their rankings in the Matbench Discovery benchmark, accessed on April 1st, 2025. The models are ranked by their highest F1 scores as follows, excluding DPA-2.4-7M: SevenNet-MF-ompa $>$ GRACE-2L-OAM $>$ Orb-v2 $>$ MatterSim-v1-5M $>$ MACE-MPA-0 $>$ SevenNet-l3i5 $>$ MACE-MP-0. This coherence further substantiates the reliability and robustness of both the LAMBench and Matbench Discovery benchmark frameworks[24] in evaluating LAMs for Inorganic Materials force field tasks. Within the Catalysis domain, DPA-2.4-7M again achieves the lowest error, significantly surpassing the other models. In the Biomolecules and Supramolecules domain, the DPA-2.4-7M model achieves the highest generalizability, distinctly outperforming all other LAMs. The remaining models show comparable performances, except for MACE-MP-0 and MACE-MPA-0, which fall noticeably behind.
Table 2 reveals that MACE-MPA-0, characterized by a larger parameter set and trained on an extensive dataset, exhibits a dimensionless error 0.05 lower than MACE-MP-0, suggesting enhanced generalizability. This trend is corroborated by the domain-wise error metrics in Figure 2 (a), which demonstrate superior performance across most domains, with the exception of Catalysis, where a minor decline is noted. The Inorganic Materials domain shows the most significant improvement, with an approximately 30 percent reduction in error. Meanwhile, SevenNet-MF-ompa delivers overall performance on par with its lighter counterpart, SevenNet-l3i5; however, its enhanced accuracy in the Inorganic Materials and Reactions domains comes at the cost of reduced generalizability elsewhere, indicating potential overfitting to the Inorganic Materials domain.
The DPA-2.4-7M model demonstrated the best overall performance across all investigated domains. This success can be attributed to its training on 31 datasets, detailed in Supporting Information LABEL:tab:si-datasets, which span multiple domains using a multitask training strategy[16, 4]. In this approach, a common feature extraction structure, the descriptor, is linked to multiple task heads. Each task head is trained on one of the 31 datasets, while the common descriptor is simultaneously trained on all datasets. Consequently, the descriptor is designed to extract the shared knowledge from all the training datasets. For a fair comparison, the task head trained with the MPtrj dataset, which serves as a subset of the training data for all LAMs, is consistently used to assess the generalization error. Thus, the DPA-2.4-7M model’s exceptional performance in the Small Molecules, Biomolecules and Supramolecules, and Catalysis domains can be attributed to the descriptor, which was trained using diverse datasets from domains such as Small Molecules (including the SPICE2 and Yang2023ab datasets) and Catalysis (including the OC20 and OC22 datasets).
It is important to note that certain test datasets are labeled with an exchange-correlation (XC) functional that differs from the one used for labeling the training datasets, such as MPtrj, for the LAMs. For instance, in the domain of Small Molecules, the ANI test set is labeled by $\omega$B97X/6-31G* functional. To further investigate the impact of XC functional mismatch between the model’s training data and the reference functional used in generalizability tests, we reassessed the generalization error of the DPA-2.4-7M model using task heads that align as closely as possible with the XC functional employed for labeling the test datasets. This alignment resulted in a substantial reduction in generalization error across several domains. Specifically, the domain error metric decreased from 0.19 to 0.09 for Small Molecules and from 0.25 to 0.17 for Biomolecules and Supramolecules (see Figure 2 (a)). These findings suggest that XC functional mismatch significantly contributes to the generalization error of the LAM model, highlighting the necessity of training models with multiple XC functional fidelities[59].
@@
The generalizability of LAMs in property calculation tasks is assessed and summarized in Table 2. Compared to force field tasks, property calculation tasks offer greater discriminative power among models. The DPA-2.4-7M model again achieves the best performance, with the lowest error of 0.208—representing nearly a two-thirds reduction in error relative to the worst-performing model. The domain-wise error metric is shown in Figure 2(b), with the detailed raw measurements available in Supporting Information Tables S-3 and S-4.
@@
In the Inorganic Materials domain, where the MDR phonon benchmark is conducted, the SevenNet-MF-ompa, MatterSim-v1-5M, and GRACE-2L-OAM models demonstrate superior performance, with MACE-MPA-0 following closely. Notably, the conservative models exhibit significantly lower errors compared to the non-conservative Orb-v2 model, which directly predicts force. Accurate phonon predictions necessitate the calculation of the force constant matrix, representing the second-order derivative of a force field. Therefore, conservativeness and smoothness are crucial for achieving precise phonon predictions, consistent with previous findings in Ref.[26]. Additionally, the performance difference between MACE-MP-0 and its more expressive counterpart, MACE-MPA-0, suggests that enhanced generalizability in force field prediction tasks can lead to improved performance in a property prediction task.
In the Small Molecules domain, where the TorsionNet500 benchmark is conducted, the dimensionless error values are notably higher compared to those in the Inorganic Materials domain. The multi-task trained LAM, DPA-2.4-7M, demonstrates the lowest error at 0.35, outperforming other models primarily trained on datasets within the Inorganic Materials domain. Utilizing the task head trained with the SPICE2 dataset, which employs a closely aligned XC as the reference DFT method, reduces the error to 0.19. Despite these improvements, performance remains significantly lower than that of domain-specific models such as MACE-OFF23[60], which achieves a dimensionless error of 0.034 on TorsionNet500 benchmark. The discrepancy between the performance of LAMs and MACE-OFF23 highlights the necessity of focusing on domains beyond Inorganic Materials to enhance generalizability in property prediction tasks. This also indicates that current LAMs remain considerably less accurate than domain-specific models and are still distant from achieving the capabilities of a universal PES model.
The adaptation of the pretrained DPA-2.4-7M model across eight Matbench property regression tasks is detailed in Table 3. The fine-tuned model consistently surpasses the model with the same architecture trained from scratch in terms of accuracy. Despite the improvements achieved by the finetuned DPA-2.4-7M model, a discrepancy remains between its performance and the results showcased on the Matbench leaderboard.
@@
Notably, a property predictor adapted from the pretrained MatterSim[61] achieves performance comparable to leading task-specific models on the Matbench leaderboard. Furthermore, JMP, pretrained on the OC20, OC22, ANI-1x, and Transition-1x datasets[16], significantly surpasses existing models, achieving state-of-the-art accuracy. This underscores the critical importance of pretraining for achieving superior model accuracy in downstream property prediction tasks and reinforces the broader vision of LAMs as versatile, high-performing surrogates capable of addressing a wide range of scientific challenges.
@@
Table 3: Adaptability Test: The accuracy of property fine-tuning of DPA-2.4-7M across eight Matbench regression tasks.
@@
Task (Unit) DPA-2.4-7M From scratch DPA-2.4-7M Finetune11footnotemark: 1 Matbench[37] leaderboard MatterSim[61] JMP[16]
MP $E_{\text{form}}$ (meV/atom) 30.1 24.1 17.0 - 10.1
MP Gap (eV) 0.256 0.206 0.156 0.129 0.091
JDFT2D (meV/atom) 43.68 34.09 33.19 32.76 29.94
Phonons (cm-1) 43.55 34.00 28.76 26.02 20.57
Dielectric (unitless) 0.418 0.300 0.271 0.252 0.249
Log KVRH ($\log_{10}\text{GPA}$) 0.061 0.052 0.049 0.049 0.045
Log GVRH ($\log_{10}\text{GPA}$) 0.079 0.068 0.067 0.061 0.059
Perovskites (eV/unitcell) 0.054 0.032 0.027 - 0.026
@@
Results were obtained using NVIDIA A800-SXM4-40GB GPUs with the maximum permissible batch size. Performance may degrade on NVIDIA L4-24GB GPUs with smaller batch sizes as illustrated in Table S-5, which are used by default in the workflow for cost management purposes.
@@
Figure 3: Distribution of inference time, normalized by the number of atoms, measured across 1,600 randomly selected configurations. Lower values indicate higher efficiency.
The stability of the LAMs is evaluated using the instability metric $M^{m}_{\mathrm{IS}}$, as presented in Table 2. The conservative models, such as MatterSim-v1-5M, exhibit minimal energy drift throughout the evaluation, indicating stable simulations. In contrast, the non-conservative model, Orb-v2, shows instability metric several orders of magnitude larger, reflecting instability in the simulation. In general, models with a greater number of trainable parameters and trained on larger datasets exhibit higher stability, as reflected in reduced energy drift. This trend is evident in the MACE, and SevenNet model families as shown in Supporting Information Table S-6.
@@ III Discussion and Conclusion
This study introduces a comprehensive benchmarking framework, designated as LAMBench, for evaluating large atomic models (LAMs). It is designed to assess the extent to which LAMs can be used as versatile, out-of-the-box tools capable of advancing scientific discovery across a broad range of contexts. We emphasize three essential requirements for realizing this vision: generalizability to atomic systems across diverse research domains, adaptability to novel tasks, and applicability in real-world simulations with regard to stability and efficiency.
@@
In this study, we benchmarked eight leading LAMs, released before April 1, 2025. DPA-2.4-7M consistently demonstrated superior performance in terms of generalizability compared to all other LAMs. In the force field generalizability test, DPA-2.4-7M was followed by GRACE-2L-OAM, SevenNet-l3i5, SevenNet-MF-ompa, MACE-MPA-0, and Orb-v2, all exhibiting remarkably similar accuracy levels. A trend of overfitting to the Inorganic Materials domain was observed in some LAMs. However, in the generalizability test for property calculation, more pronounced differences between the models are evident. In the phonon calculation test, the disparities are mainly attributed to the conservativeness of the models, with conservative models significantly outperforming the non-conservative ones. In the torsion profile calculation task within the Small Molecules domain, where most LAMs, except DPA-2.4-7M and SevenNet-l3i5, exhibit a dimensionless error of $\geq 0.5$. Although the XC-adapted DPA-2.4-7M model achieves the lowest generalization error of 0.19, it remains significantly higher than MACE-OFF23, a model specifically trained for the Small Molecules domain, which has an error of 0.034.
The LAMBench results highlight a substantial gap between current LAMs and the ideal universal model, envisioned as an out-of-the-box simulation tool for scientific discovery. This gap is primarily due to limited generalizability beyond the Inorganic Materials domain, which is largely attributed to training datasets predominantly sourced from this domain. While DPA-2.4-7M employs a multi-task training scheme, simultaneously utilizing data from multiple domains to enhance generalization capabilities, its accuracy remains inferior to models specifically oriented towards individual domains. To bridge this gap, future research might focus on developing advanced model architectures for improved generalizability, more efficient training methods to learn shared knowledge across diverse domains, and a balanced distribution of training data across various research fields. These directions hold promise for advancing the capabilities of LAMs in scientific discovery.
The capability to adjust model fidelity is proposed as an essential feature for LAMs, given that different research domains necessitate varying levels of accuracy in DFT calculations. For instance, the Inorganic Materials domain typically requires GGA functionals, whereas the Small Molecules domain often demands hybrid functionals. Notably, the force field error in the Small Molecules domain for DPA-2.4-7M decreases from 0.19 to 0.09 when switching from PBE/PBE+U to the $\omega$B97X-D functional. This underscores the importance of adaptable fidelity in achieving optimal performance across diverse research applications.
Regarding the efficiency of using LAMs in simulations, the non-conservative model Orb-v2 demonstrates significantly higher performance compared to the conservative models. Among the conservative models, the fastest model is 7.7 times quicker than the slowest, a disparity that is notably larger than the variations observed in generalizability. This observation underscores the importance of considering the efficiency in the development of LAMs, particularly for speed-critical applications such as long-time molecular dynamics simulations.
@@
In this study, adaptability is exclusively benchmarked for the DPA-v2.4-7M model, as most implementations of LAMs are primarily confined to force field prediction and are not readily adaptable to property prediction tasks. The adaptability test for DPA-2.4-7M demonstrates that a pretrained model can offer substantial advantages in downstream property prediction tasks compared to models trained from scratch, particularly in data-scarce scenarios. Despite DPA-2.4-7M’s strong performance in generalizability tests, a noticeable gap remains between its accuracy and those of state-of-the-art property prediction models. This highlights a significant opportunity for further enhancing the adaptability of the DPA-2.4-7M model representing a crucial step towards the development of ideal universal models. Furthermore, it is advisable for LAM developers to enhance their models with the capability to adapt for property prediction. This advancement could open up new avenues for their application in scientific discovery.
In the future, LAMBench should be enhanced to more accurately reflect the LAMs’ performance in real-world applications as an out-of-box simulation tool. This enhancement requires the collection of additional test datasets for the force field generalizability test and the incorporation of more property calculation workflows in the property generalizability test. It is important to note that the design of property calculation tasks requires careful consideration. Directly borrowing all property calculations from applications may not be optimal, as these calculations often necessitate long time-scale and large spatial scale MD simulations, such as for ionic diffusion constants, which demand substantial computational resources. Ideally, property benchmarks should be designed as representative examples that reflect performance in the calculation of domain-specific properties while remaining computationally feasible.

### v2

@@ II.2 Benchmark results
Table 2 presents the LAMBench Leaderboard, showcasing the performance of LAMs in terms of their generalizability and applicability. Currently, adaptability is excluded from the comparison due to the absence of this test for most of the LAMs under evaluation. This leaderboard aims to strike a balanced comparison between key performance aspects and is intended to offer insights into model suitability for applications relevant to real-world scientific discovery. To reduce discrepancies in the comparisons between multi-task and single-task trained models, we employed the MPtrj task head for the DPA-3.1-3M and DPA-2.4-7M models, and the MPA modality for the SevenNet-MF-ompa model, unless stated otherwise.
@@
DPA-3.1-3M 0.175 0.322 0.261 0.572
Orb-v3 0.215 0.414 0.396 0.000
DPA-2.4-7M 0.241 0.342 0.617 0.039
GRACE-2L-OAM 0.251 0.404 0.639 0.309
Orb-v2 0.253 0.601 1.341 2.649
SevenNet-MF-ompa 0.255 0.455 0.084 0.000
MatterSim-v1-5M 0.283 0.467 0.393 0.000
MACE-MPA-0 0.308 0.425 0.293 0.000
SevenNet-l3i5 0.326 0.397 0.272 0.036
MACE-MP-0 0.351 0.472 0.296 0.089
@@
Figure 2: Dimensionless error metrics for generalizability tasks across different domains. (a) Dimensionless error metrics for force field tasks across three distinct domains. (b) Dimensionless error metrics for property-calculation tasks. To reassess DPA-3.1-3M with enhanced domain alignment, referred to as DPA-3.1-3M-DomainMatch, we employed task heads trained on SPICE2 for the Molecules domain, OMAT24 for the Inorganic Materials domain, and OC20M for the Catalysis domain.
In the force field generalizability task, as illustrated in Table 2, performance is evaluated using the dimensionless error $\bar{M}_{\mathrm{FF}}^{m}$. Notably, DPA-3.1-3M demonstrates substantially greater generalizability compared to other LAMs.
@@
To further investigate the generalizability of these LAMs across specific domains, we present the domain-wise dimensionless error metric in Figure 2(a). In general, most LAMs exhibit relatively low inference errors in the domains of Molecules and Inorganic Materials, while showing inferior performance in the domain of Catalysis. For Molecules, DPA-3.1-3M ranks first, closely followed by DPA-2.4-7M and Orb-v3, and outperforms all other models in this domain. In the Inorganic Materials domain, the comparative performance of the models shows a similar pattern to their rankings in the Matbench Discovery benchmark, accessed on August 1st, 2025. These models, excluding DPA2-2.4-7M, are ranked by their lowest MAE in the following order: SevenNet-MF-ompa << GRACE-2L-OAM << DPA-3.1-3M << Orb-v3 << MatterSim-v1-5M << MACE-MPA-0 << Orb-v2 << SevenNet-l3i5 << MACE-MP-0. This coherence further substantiates the reliability and robustness of both the LAMBench and Matbench Discovery benchmark frameworks[25] in evaluating LAMs for Inorganic Materials force field tasks. Within the Catalysis domain, DPA-3.1-3M again achieves the lowest error, significantly surpassing the other models.
Table 2 indicates that MACE-MPA-0, which features an expanded set of parameters and is trained on a more comprehensive dataset, achieves a dimensionless error of 0.308. It is 0.043 lower than the error of MACE-MP-0, which stands at 0.351, suggesting improved generalizability. This finding is further supported by the domain-specific error metrics presented in Figure 2(a), where MACE-MPA-0 demonstrates superior performance across all three domains. Notably, in the Inorganic Materials domain, there is an approximate 25% reduction in error. Similarly, the SevenNet-MF-ompa model shows a comparable improvement over its lighter version, SevenNet-l3i5.
The DPA-3.1-3M model[61] demonstrates the best overall performance across all investigated domains. This achievement is likely due to its multi-task training strategy[16, 4, 17] that train the model on 31 datasets spanning multiple domains, detailed in Supporting Information LABEL:tab:si-datasets. The model’s outstanding performance in the Molecules and Catalysis domains can be attributed to the descriptor, which was trained using diverse datasets from domains such as Molecules (including the SPICE2 and Yang2023ab datasets) and Catalysis (including the OC20 and OC22 datasets). The DPA-2.4-7M model, trained using the same multitask strategy, ranks third, showing a noticeable gap from DPA-3.1-3M and Orb-v3, largely due to its limited capacity.
@@
The generalizability of LAMs in property calculation tasks is assessed and summarized in Table 2. Compared to force field tasks, property calculation tasks offer greater discriminative power among models. The DPA-3.1-3M model again achieves the best performance, with the lowest error of 0.322—representing nearly a fifty percent reduction in error relative to the worst-performing model. The domain-wise error metrics are presented in Figure 2(b), with the detailed raw measurements available in Supporting Information Tables S-3, S-4, S-5, S-6 and S-7.
@@
In the Inorganic Materials domain, where the MDR phonon and elasticity benchmarks are conducted, the conservative models exhibit significantly lower errors compared to the non-conservative Orb-v2 model, which directly predicts force. Accurate phonon and elasticity predictions necessitate the calculation of the force constant matrix, representing the second-order derivative of a force field. Therefore, conservativeness and smoothness are crucial for achieving precise predictions, consistent with previous findings in Ref.[27]. Additionally, the performance difference between MACE-MP-0 (0.14) and its more expressive counterpart, MACE-MPA-0 (0.06), suggests that enhanced generalizability in force field prediction tasks can lead to improved performance in a property prediction task. In both benchmarks, the Omat24 task head of DPA-3.1-3M, denoted as DPA-3.1-3M-DomainMatch, demonstrates improved performance over the MPtrj task head, underscoring the benefits of more extensive training data.
[…]
In the Catalysis domain, where the OC20NEB-OOD benchmark is conducted, all LAMs exhibit limited generalizability. The top-performing model, DPA-3.1-3M-DomainMatch, utilizing the OC20M task head, still presents a relatively high dimensionless error of 0.53. In contrast, domain-specific models such as EquiformerV2-31M-S2EF-OC20-All+MD[64, 65] and eSCN-L6-M3-Lay20-S2EF-OC20-All+MD[66, 65] achieve substantially lower dimensionless errors of 0.31 and 0.33, respectively. As illustrated in Table S-7, LAMs achieve notably lower errors when predicting the reaction energy ($dE$) as opposed to the reaction barrier ($E_{a}$). This discrepancy indicates that improving the performance of LAMs necessitates increased focus on the transition states.
@@
The adaptation of the pretrained DPA-3.1-3M and DPA-2.4-7M models across eight Matbench property regression tasks is detailed in Table 3. The fine-tuned model consistently surpasses the model with the same architecture trained from scratch in terms of accuracy. Notably, a property predictor adapted from the pretrained MatterSim[67] achieves performance comparable to leading task-specific models on the Matbench leaderboard. Furthermore, JMP, pretrained on the OC20, OC22, ANI-1x, and Transition-1x datasets[16], significantly surpasses existing models, achieving state-of-the-art accuracy. Although a performance gap remains between DPA-3.1-3M and the state-of-the-art (SOTA) results, its superior performance compared to DPA-2.4-7M demonstrates that better generalizability in force field prediction tasks also translates to improved adaptability in property fine-tuning tasks. This underscores the critical importance of pretraining for achieving superior model accuracy in downstream property prediction tasks and reinforces the broader vision of LAMs as versatile, high-performing surrogates capable of addressing a wide range of scientific challenges.
@@
Table 3: Adaptability Test: The accuracy of property fine-tuning of DPA models across eight Matbench regression tasks. 11footnotemark: 1
@@
Task (Unit) DPA-2.4-7M From scratch DPA-2.4-7M Finetune DPA-3.1-3M From scratch DPA-3.1-3M Finetune Matbench[56] leaderboard MatterSim[67] JMP[16]
MP $E_{\text{form}}$ (meV/atom) 31.1 24.7 24.2 13.9 17.0 - 10.1
MP Gap (eV) 0.285 0.261 0.255 0.147 0.156 0.129 0.091
JDFT2D (meV/atom) 46.93 32.23 41.80 31.29 33.19 32.76 29.94
Phonons (cm-1) 41.10 29.05 41.41 25.13 28.76 26.02 20.57
Dielectric (unitless) 0.431 0.292 0.438 0.297 0.271 0.252 0.249
Log KVRH ($\log_{10}\text{GPA}$) 0.062 0.052 0.060 0.051 0.049 0.049 0.045
Log GVRH ($\log_{10}\text{GPA}$) 0.079 0.066 0.079 0.061 0.067 0.061 0.059
Perovskites (eV/unitcell) 0.061 0.047 0.043 0.035 0.027 - 0.026
@@
Results were obtained using 4 * NVIDIA L4-24GB GPUs with the maximum permissible batch size. For the MP $E_{\text{form}}$ and MP Gap tasks, the reported accuracy may not reflect the fully converged results due to insufficient training epochs.
@@
Figure 3: Distribution of inference time, normalized by the number of atoms, measured across 900 randomly selected configurations. Lower values indicate higher efficiency.
The stability of the LAMs is evaluated using the instability metric $M^{m}_{\mathrm{IS}}$, as presented in Table 2. The conservative models, such as Orb-v3 , exhibit minimal energy drift throughout the evaluation, indicating stable simulations. In contrast, the non-conservative model, Orb-v2, shows instability metric several orders of magnitude larger, reflecting instability in the simulation. The relatively high instability metric observed in DPA-3.1-3M is due to a failed simulation, potentially caused by the relatively low training weight assigned to the MPtrj task head, as detailed in LABEL:tab:si-datasets. Notably, utilizing the Omat24 task head reduces the instability metric to zero. In general, models with a greater number of trainable parameters and trained on larger datasets exhibit higher stability, as reflected in reduced energy drift. This trend is evident in the MACE, and SevenNet model families as shown in Supporting Information Table S-8.
@@ III Discussion and Conclusion
This study introduces a comprehensive benchmarking framework, designated as LAMBench, for evaluating large atomistic models (LAMs). It is designed to assess the extent to which LAMs can be used as versatile, out-of-the-box tools capable of advancing scientific discovery across a broad range of contexts. We emphasize three essential requirements for realizing this vision: generalizability to atomistic systems across diverse research domains, adaptability to novel tasks, and applicability in real-world simulations with regard to stability and efficiency.
@@
The LAMBench results highlight a substantial gap between current LAMs and the ideal universal model, envisioned as an out-of-the-box simulation tool for scientific discovery. This gap is primarily due to limited generalizability beyond the Inorganic Materials domain, which is largely attributed to training datasets predominantly sourced from this domain. More specialized datasets are needed to improve cross-domain generalization, such as those containing transition states in the Catalysis domain. Although DPA-3.1-3M employs a multi-task training scheme, utilizing data from multiple domains to enhance its generalization capabilities and outperforming other single-task trained models in most situations, its accuracy remains inferior to models specifically tailored for individual domains. To bridge this gap, future research might focus on developing advanced model architectures for improved generalizability, more efficient training methods to learn shared knowledge across diverse domains, and a balanced distribution of training data across various research fields. These directions hold promise for advancing the capabilities of LAMs in scientific discovery.
The capability to adjust model fidelity is proposed as an essential feature for LAMs, given that different research domains necessitate varying levels of accuracy in DFT calculations. For example, the Inorganic Materials domain generally necessitates GGA functionals, whereas the Molecules domain often requires hybrid or higher-level functionals. Training models across various XC functional fidelities offers considerable flexibility for downstream applications and should thus be actively promoted within the community [68].
Regarding the efficiency of using LAMs in simulations, the non-conservative model Orb-v2 demonstrates significantly higher performance compared to the conservative models. Among the conservative models, the fastest model is 7.6 times quicker than the slowest, a disparity that is notably larger than the variations observed in generalizability. This observation underscores the importance of considering the efficiency in the development of LAMs, particularly for speed-critical applications such as long-time molecular dynamics simulations.
@@
In this study, adaptability is exclusively benchmarked for the DPA models, as most implementations of LAMs are primarily confined to force field prediction and are not readily adaptable to property prediction tasks. The adaptability test for DPA-3.1-3M and DPA-2.4-7M demonstrates that a pretrained model can offer substantial advantages in downstream property prediction tasks compared to models trained from scratch, particularly in data-scarce scenarios. The comparison between these two models further suggests that greater generalizability in force field prediction tasks translates to improved adaptability in property fine-tuning tasks. Additional research is needed to validate this correlation across other LAMs. Furthermore, it is advisable for LAM developers to enhance their models with the capability to adapt for property prediction. This advancement could open up new avenues for their application in scientific discovery.
Currently, LAMBench only includes twelve OOD datasets for force field prediction and five for property calculation across three research domains, supporting eight adaptability and two applicability tests. In the future, LAMBench should incorporate more comprehensive downstream test cases, encompassing both force field and property calculation scenarios, to more accurately reflect the performance of LAMs in real-world applications as an out-of-the-box simulation tool. This enhancement necessitates dynamically adjusting the force field generalizability test datasets to accommodate emerging trends in training dataset development, which may render some OOD generalizability test cases in-distribution. Consequently, it will be essential to replace these test cases with more challenging alternatives. It is important to note that the design of property calculation tasks requires careful consideration. Directly borrowing all property calculations from applications may not be optimal, as these calculations often necessitate long time-scale and large spatial scale MD simulations, such as for ionic diffusion constants, which demand substantial computational resources. Ideally, property benchmarks should be designed as representative examples that reflect performance in the calculation of domain-specific properties while remaining computationally feasible.